Multiply and solve for y are the instructions.
y + 3 = 15
Please show step by step in solving the equation.
Please e-mail asap. Thank you. JC

Answer by brandonpark2889(31)

Are you sure it is y+3=15? If it is, then:
y+3=15
-3 -3
======
y=12.
Simple
---
I love math, period.